OpenStudy (anonymous):

Factor completely. 4x^2-4x+2xy-2y

OpenStudy (anonymous):

thats better!

OpenStudy (anonymous):

(4x+2y)(x-1)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

= 4x^2 + 2xy -4x -2y = 2x(2x + y) -2(2x +y)) = (2x-2)(2x+y) = 2(x-1)(2x+y)

OpenStudy (anonymous):

where's my medal pls ?? :)